# FLARE: Faithful Logic-Aided Reasoning and Exploration

### v4

@@ 5.3 Faithful Reasoning Improves Performance
Table 5: The table shows how the percentage of unique emergent inferences in search, overlapping relations between code and search, and unused relations in code impact answer correctness.
As described in Section 3, using FLARE allows us to measure the faithfulness of the LLM reasoning process by comparing the simulated problem space traversals ${\Phi}^{\prime}_{\text{\emph{gen}}}$ with actual traces ${\Phi}^{\prime}$ produced from a symbolic Prolog solver. To do this, we initially compute the percentage of syntactically correct executable code each LLM produces. We can see from the right part of Figure 3 that all of the models are capable of producing correct executable Prolog code in $67\%$ of cases on average and $\geq 50\%$ of cases at the very least. This shows that the simulated searches ${\Phi}^{\prime}_{\text{\emph{gen}}}$ can be considered a representative sample that will be further used to accurately measure the faithfulness of the simulated search w.r.t. the generated code. After measuring the reasoning faithfulness for each model, we want to understand what impact it has on the performance of the LLM. In Figure 2, we segment the models w.r.t. their ROUGE-Lsum scores. The results show that model performance is strongly positively correlated with reasoning faithfulness. However, we also observe in the left part of Figure 3 that executing semantically precise code results in an accurate answer only in $47\%$ of cases on average. Indeed, having a simulated search trace with a ROUGE-Lsum faithfulness score of $1$, would be equivalent to simply executing the program as proposed in F-CoT. Yet we have priorly shown that F-CoT struggles with reasoning tasks that are hard to formalise and require multi-hop commonsense and soft reasoning. These two discoveries show that optimal LLM reasoning, conditioned on a search in the problem space, should be increasingly faithful toward the facts, relations and the search strategy defined within the code while simultaneously maintaining the capability for soft-reasoning along more abstractly defined concepts. Our results show that FLARE allows LLMs to maintain a similar reasoning capacity.
5.4 What is important during the search?
Model Avg. hops per Paths Hallucination (%) Unutilised knowledge (%) Llama-3.1-8B 9.4 63.3 62.9 CmDR 6.7 54.7 56.9 CmDR+ 7.2 54.3 56.3 GPT-3.5 5.5 49.3 52.1
Table 6: The table shows the changes in simulated search statistics when using FLARE w.r.t model scale from 8B to 100B+. Hallucinations refer to facts and predicates only used in trace, while unutilised knowledge relates to the facts and relations only seen in the code.
We expand the analysis of the simulated search traces to detect the reasons which can lead to optimal reasoning within an LLM. For this purpose, we calculate several statistics, like the average number of explored paths, average and total hops and failures per path, for each model during the simulated traversal. The failure in a path is an invalidation of a solution for a sub-goal explored during the search, which is used for backtracking, as explained in Section 3. Calculating these statistics is simple as the search component of FLARE, seen in Figure 1, is a structured simulation of a Prolog trace, where each line contains a hop of reasoning inference. We split these statistics for the reasoning paths that lead to correct or incorrect outcomes. Our results in Table 4 show that LLM performance and reasoning optimally are not directly connected to the amount of explored paths or multi-hop inferences per path. We also see that traces that lead to incorrect answers have a higher number of failures per path and in total. We explain this phenomenon with the hypothesis that LLMs with traces that were optimal for reasoning and led to correct answers could skip exploring degenerate solutions due to strong commonsense reasoning capabilities. Further analyses focus on identifying inconsistencies and failure modes (Section 3.2). By comparing relations in code with those in search traces, we measure emergent hallucinations and unused relations, highlighting areas of sub-optimal reasoning. Additionally, we assess the uniqueness of emergent facts per inference hop, which indicates the extent of problem-space exploration (Table 5). The results in Table 5 show consistently over each model that, on average, traces that lead to correct answers had a higher percentage of unique emergent facts (UEF) and overlap in the relations (OR) used between the code and search, while the portion of underutilized relations (UR) was lower. This means that optimal reasoning with an LLM requires a great degree of problem-space exploration with fewer relation hallucinations during the search and more relation utilization from the defined code. This aligns with our prior discoveries, which show a strong correlation between simulated search faithfulness towards the formalised code and model performance. Our framework FLARE has these reasoning patterns ingrained within its inference pipeline.
@@ 5.4 What is important during the search?
5.5 The effect of scale
We want to assess the impact of the number of parameters in the model on the overall performance and faithfulness. The results in Figure 4 show no precise relation between model scale, performance and faithfulness. However, scaled models from the same family, i.e. CmDR (30B) and CmDR+ (100B), show improvements in reasoning faithfulness and model performance. We can also see in Table 6 that as the model size increases, the average number of hops and the portion of hallucinations and unutilised knowledge decreases. This further confirms our prior assumptions that models with strong commonsense soft-reasoning capabilities can skip steps during the search while maintaining the knowledge and structure of the traversal strategy outlined in the code.
6 Conclusion
This work introduces FLARE, a novel approach for logic-aided interpretable formalisation and reasoning with simulated search over the problem space. We show that models of varying scales obtain state-of-the-art results compared to prompting paradigms like CoT and F-CoT. We further pinpoint that using FLARE allows us to perform soft-reasoning with simulated search, making it flexible for diverse reasoning benchmarks. We introduce a method to measure model reasoning faithfulness w.r.t. the problem formalization ingrained within FLARE. Our results show that model performance is positively correlated with the faithfulness of the reasoning process. The systematic studies of the method show the benefits of using simulated search compared to natural language reasoning and external symbolic solvers. We further show that using FLARE allows us to interpretably and rigorously detect hallucinations and sub-optimal and inconsistent reasoning patterns.
Reproducibility Report
To reproduce the results of our study, we provide the complete codebase, processing pipelines and prompts for each dataset. The only model hyper-parameter we explicitly fix is the temperature for greedy decoding. We also make the inference of all of the models using FLARE, F-CoT and CoT across all of the datasets publicly available for further experimentation and exploration.
Appendix A Appendix
A.1 LLM Prompts
We define straight-forward prompts for generating plan, code and search simulation in FLARE, which can be observed in Table 7.
A.2 Dataset Statistics
@@
Task Prompt Description Plan Generation Generate an explanation and analysis, and plan to generate a prompt for writing a swi-prolog code for the last task. The 3 sections should be exactly outlined. Your plan should show enough intermediate reasoning steps towards the answer. Construct the plan as much as you can and describe the logic specifically. When constructing the plan for the code prompt, actively use swi prolog search capabilities. Detailed instructions for generating an outline and plan, with an emphasis on reasoning steps and using Prolog’s search capabilities. Code Generation Write a Prolog code to solve using the plan. If there are unknown or stochastic atoms or predicates, fill in the values for them as a logical assumption and add a comment in the same line Assumed atom/predicate”. Do not use write and read commands within the code. The code should be very detailed and utilize swi prolog capabilities to the fullest. To run the program, at the end create a predicate named ”query” that returns the correct numerical answer. The last line of the program should be the commented-out driver predicate ”query”. Write only the code. Instructions for generating a Prolog code based on the plan with assumptions for unknown atoms. Emphasizes code details and a final “query” predicate. Simulated Search Ignoring the read commands, explicitly write out the search paths that are explored by the code: #### Here are the paths [Starting Search Simulation]: #### [Path 1]: A task to simulate and display the search paths that the Prolog code would follow during execution. Final Answer Given the plan, the code and the explored search paths answer the question above. Answer with the correct numerical answer. ##### Here is the answer: Final prompt asking for the correct numerical answer based on the previous steps.
A.3 FLARE Pseudo-code
Below, we present the pseudo-code for the execution of the plan, code, and search procedures in FLARE. The pseudo-code describes the modular pipeline in FLARE for tackling natural language queries with faithful simulated search.
  • Plan Generation: This stage creates a structured natural language outline of the reasoning process, breaking down the query into logical steps and analysis. The plan serves as the foundation for formalization into a logic-based representation.
  • Code Generation: Based on the generated plan, a logic programming code (e.g., in Prolog) is synthesized. This code formalizes the query into a set of facts, relations, and goals, which collectively define the problem space for reasoning.
  • Search Simulation: The generated code is utilized to simulate a search trace over the problem space. This includes iterative reasoning, backtracking when goals are unmet, and extracting emergent facts or relations during the process.
Each of these stages is implemented as a modular component. The generation from each of the stages feeds into the next, allowing seamless integration and incremental improvement in reasoning accuracy. A detailed pseudo-code is provided below in Algorithm 1.
@@ A.3 FLARE Pseudo-code
Question ”Can Reiki be stored in a bottle?”
@@
Code % Define the nature of Reiki energy reiki_energy(spiritual_energy). reiki_energy(channeling_through_touch). % Define the feasibility of storing energy in a bottle store_energy_in_bottle(feasible) :- reiki_energy(spiritual_energy), reiki_energy(channeling_through_touch). % Define the assumption that storing Reiki energy in a bottle is not a common practice not_common_practice(store_reiki_in_bottle). % Query to determine if Reiki can be stored in a bottle query :- store_energy_in_bottle(feasible), not_common_practice(store_reiki_in_bottle).
Search Paths Path 1: 0: Start of execution: Beginning Search 1: Call: query 2: Call: store_energy_in_bottle(feasible) 3: Call: reiki_energy(spiritual_energy) 4: Call: reiki_energy(channeling_through_touch) 5: Call: not_common_practice(store_reiki_in_bottle) 6: Call: feasible=no 7: Fail: feasible=no — ’Result’: ’Search Failed’ 8: Fail: query — ’Result’: ’Search Failed’
Answer ”No, Reiki energy cannot be stored in a bottle based on the logical evaluation of its abstract, non-physical nature and traditional practices of Reiki.”
@@
Domain Dataset Shots Test Samples Example Math Word Problems GSM8K 8 1,319 Q: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? A: 3 SVAMP 8 1,000 Q: Dan had \$3 left with him after he bought a candy bar. If he had \$4 at the start, how much did the candy bar cost?A: 1 MultiArith 8 600 Q: A pet store had 13 siamese cats and 5 house cats. During a sale they sold 10 cats. How many cats do they have left? A: 8 ASDiv 8 2,096 Q: Adam has five more apples than Jackie. Jackie has nine apples. How many apples does Adam have? A: 14 AQuA 8 254 Q: A man walks at 5 kmph for 6 hrs and at 4 kmph for 12 hrs. His average speed is Answer option: A)4 1/3 km/h, B)7 2/3 km/h, C)9 ½ km/h, D)8 km/h, E)81 km/h A: A Multi- hop QA StrategyQA 6 2,290 Q: Did Aristotle use a laptop? A: False Date Understanding 10 359 Q: Yesterday was April 30, 2021. What is the date tomorrow in MM/DD/YYYY? A: ”05/02/2021” Sports Understanding 10 977 Q: Is the following sentence plausible? Lionel Messi was called for icing? A: False Relational Inference CLUTRR 8 1,042 Q: [Carlos] is [Clarence]’s brother. [Carlos] and his sister, [Annie], went shopping. asked her mom [Valerie] if she wanted anything, but [Valerie] said no. How is [Valerie] related to [Clarence]? A: ”mother”

### v5

@@ Acknowledgments
Appendix A Appendix
Figure 3: The figure shows the percentage of executable code per model (top) and the accuracy of the executable code when answering the queries (bottom).
@@ Appendix A Appendix
A.1 The effect of scale
Model Avg. hops per path Hal. (%) UK. (%)
Llama-3.1-8B 9.4 63.3 62.9
CmDR 6.7 54.7 56.9
CmDR+ 7.2 54.3 56.3
GPT-3.5 5.5 49.3 52.1
Table 6: Changes in simulated search statistics when using FLARE across model scales (8B to 100B+). Hallucinations (Hal.) are facts/predicates used only in the trace, while unutilised knowledge (UK.) denotes facts/relations appearing only in the code.
We want to assess the impact of the number of parameters in the model on the overall performance and faithfulness. The results in fig.˜4 show no precise relation between model scale, performance and faithfulness. However, scaled models from the same family, i.e. CmDR (30B) and CmDR+ (100B), show improvements in reasoning faithfulness and model performance. We can also see in table˜6 that as the model size increases, the average number of hops and the portion of hallucinations and unutilised knowledge decreases. This further confirms our prior assumptions that models with strong commonsense soft-reasoning capabilities can skip steps during the search while maintaining the knowledge and structure of the traversal strategy outlined in the code.
A.2 LLM Prompts
We define straight-forward prompts for generating plan, code and search simulation in FLARE, which can be observed in table˜8.
A.3 Dataset Statistics
@@
A.4 FLARE Pseudo-code
Below, we present the pseudo-code for the execution of the plan, code, and search procedures in FLARE. The pseudo-code describes the modular pipeline in FLARE for tackling natural language queries with faithful simulated search. (i) Plan Generation: This stage creates a structured natural language outline of the reasoning process, breaking down the query into logical steps and analysis. The plan serves as the foundation for formalization into a logic-based representation. (ii) Code Generation: Based on the generated plan, a logic programming code (e.g., in Prolog) is synthesized. This code formalizes the query into a set of facts, relations, and goals, which collectively define the problem space for reasoning. (iii) Search Simulation: The generated code is utilized to simulate a search trace over the problem space. This includes iterative reasoning, backtracking when goals are unmet, and extracting emergent facts or relations during the process. Each of these stages is implemented as a modular component. The generation from each of the stages feeds into the next, allowing seamless integration and incremental improvement in reasoning accuracy. A detailed pseudo-code is provided below in algorithm˜1.
Domain Dataset Shots Test Samples Example Math Word Problems GSM8K 8 1,319 Q: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? A: 3 SVAMP 8 1,000 Q: Dan had \$3 left with him after he bought a candy bar. If he had \$4 at the start, how much did the candy bar cost?A: 1 MultiArith 8 600 Q: A pet store had 13 siamese cats and 5 house cats. During a sale they sold 10 cats. How many cats do they have left? A: 8 ASDiv 8 2,096 Q: Adam has five more apples than Jackie. Jackie has nine apples. How many apples does Adam have? A: 14 AQuA 8 254 Q: A man walks at 5 kmph for 6 hrs and at 4 kmph for 12 hrs. His average speed is Answer option: A)4 1/3 km/h, B)7 2/3 km/h, C)9 ½ km/h, D)8 km/h, E)81 km/h A: A Multi- hop QA StrategyQA 6 2,290 Q: Did Aristotle use a laptop? A: False Date Understanding 10 359 Q: Yesterday was April 30, 2021. What is the date tomorrow in MM/DD/YYYY? A: "05/02/2021" Sports Understanding 10 977 Q: Is the following sentence plausible? Lionel Messi was called for icing? A: False Relational Inference CLUTRR 8 1,042 Q: [Carlos] is [Clarence]’s brother. [Carlos] and his sister, [Annie], went shopping. asked her mom [Valerie] if she wanted anything, but [Valerie] said no. How is [Valerie] related to [Clarence]? A: "mother"
A.5 Benefits of Prolog
Prolog is a symbolic logic-programming engine (Bowen, 1979) used for heuristic search over Horn Clauses (Chandra and Harel, 1985). It is a declarative programming paradigm (Lloyd, 1994), meaning that the code is expressed as the logic of computation. In particular, this logic is formalised as a set of facts $\mathcal{F}$ and relations $\mathcal{R}$ forming our problem space, while the final goal $\mathcal{G}$ is a first-order logic combination of them. As a default, Prolog uses a depth-first search (DFS) strategy (Bowen, 1979) for sub-goal decomposition and feasible traversal of the problem space that satisfies the goal $\mathcal{G}$. Such a traversal is referred to as the trace. At each trace step, the program can either confirm or invalidate the sub-goal using the feasibility of fact and relation combinations, expand the search tree or retry satisfying a failed sub-goal with new combinations. An example of such a search can be observed in fig.˜1. It is possible to complete an exhaustive search, exploring all possible paths that do or do not satisfy the goal. These characteristics are beneficial as we can explicitly access and segment the facts and relations that form the problem space and the search strategy used for query formalisation. As Prolog is declarative, it is sufficient to use a regexp heuristic for the segmentation, which is referred to as EXTRACT in eq.˜2 and eq.˜3. Furthermore, including exhaustive traversal traces in-context allows the LLM to simulate sub-goal decomposition, backtracking, intermediate goal invalidation, etc. We discuss this in more depth in the next paragraph.
@@ A.5 Benefits of Prolog
Task Prompt Description Plan Generation Generate an explanation and analysis, and plan to generate a prompt for writing a swi-prolog code for the last task. The 3 sections should be exactly outlined. Your plan should show enough intermediate reasoning steps towards the answer. Construct the plan as much as you can and describe the logic specifically. When constructing the plan for the code prompt, actively use swi prolog search capabilities. Detailed instructions for generating an outline and plan, with an emphasis on reasoning steps and using Prolog’s search capabilities. Code Generation Write a Prolog code to solve using the plan. If there are unknown or stochastic atoms or predicates, fill in the values for them as a logical assumption and add a comment in the same line Assumed atom/predicate". Do not use write and read commands within the code. The code should be very detailed and utilize swi prolog capabilities to the fullest. To run the program, at the end create a predicate named "query" that returns the correct numerical answer. The last line of the program should be the commented-out driver predicate "query". Write only the code. Instructions for generating a Prolog code based on the plan with assumptions for unknown atoms. Emphasizes code details and a final “query” predicate. Simulated Search Ignoring the read commands, explicitly write out the search paths that are explored by the code: #### Here are the paths [Starting Search Simulation]: #### [Path 1]: A task to simulate and display the search paths that the Prolog code would follow during execution. Final Answer Given the plan, the code and the explored search paths answer the question above. Answer with the correct numerical answer. ##### Here is the answer: Final prompt asking for the correct numerical answer based on the previous steps.
@@
Question "Can Reiki be stored in a bottle?"
@@
Code
        % Define the nature of Reiki energy
        reiki_energy(spiritual_energy).
        reiki_energy(channeling_through_touch).
        % Define the feasibility of storing energy in a bottle
        store_energy_in_bottle(feasible) :-
            reiki_energy(spiritual_energy),
            reiki_energy(channeling_through_touch).
        % Storing Reiki energy is not a common practice
        not_common_practice(store_reiki_in_bottle).
        % Query to determine if Reiki can be stored in a bottle
        query :-
            store_energy_in_bottle(feasible),
            not_common_practice(store_reiki_in_bottle).
        % :- query.
Search Paths 0: Start of execution: Beginning Search 1: Call: query 2: Call: store_energy_in_bottle(feasible) 3: Call: reiki_energy(spiritual_energy) 4: Call: reiki_energy(channeling_through_touch) 5: Call: not_common_practice(store_reiki_in_bottle) 6: Call: feasible=no 7: Fail: feasible=no | ’Result’: ’Search Failed’ 8: Fail: query | ’Result’: ’Search Failed’
Answer "No, Reiki energy cannot be stored in a bottle based on the logical evaluation of its abstract, non-physical nature and traditional practices of Reiki."